1)    A stone is thrown downward from the edge of a building 19.6 m high with an initial velocity of 4.9 m/s. Determine:

a)    The velocity of the stone just before its strikes the ground

b)    The time the stone traveled downward



1
Expert's answer
2021-05-03T10:34:27-0400

a) The velocity of the stone just before its strikes the ground is


"v=\sqrt{v_0^2+2gh}=\sqrt{4.9^2+2·9.8·19.6}=20.2\text{ m/s}."

b) The time the stone traveled downward is


"t=\frac{v-v_0}{g}=1.56\text{ s}."